OpenStudy (anonymous):

Factor 4n 2 - 5n - 6. (4n - 3)(n + 2) (4n + 3)(n - 2) (4n + 2)(n - 3)

OpenStudy (anonymous):

What do you think the correct answer is? :)

OpenStudy (anonymous):

you again lol

OpenStudy (anonymous):

:) What is your understanding of this question?

OpenStudy (anonymous):

i think it is the first one

OpenStudy (anonymous):

Okay. Let's see if you're correct:

OpenStudy (anonymous):

Trying to factor by splitting the middle term Factoring 4n^2-5n-6 The first term is, 4n2 its coefficient is 4 . The middle term is, -5n its coefficient is -5 . The last term, "the constant", is -6 Step-1 : Multiply the coefficient of the first term by the constant 4 • -6 = -24

OpenStudy (anonymous):

Is this clear to you so far?

OpenStudy (anonymous):

so so

OpenStudy (anonymous):

What isn't clear? :) I'll be happy to explain it to you.

OpenStudy (anonymous):

ok so when you multiply the coefficient of the first term do i muliply 4 tiimes neg 6 and 4 times - 24 or just six

OpenStudy (anonymous):

4 times negative 6.

OpenStudy (anonymous):

We are trying to split the middle term because this is a trinomial. In trinomials, we need to do that so we can factor by grouping. Understand? :)

OpenStudy (anonymous):

-24

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Great job! Remember that it will be -24x

OpenStudy (anonymous):

Step-2 : Find two factors of -24 whose sum equals the coefficient of the middle term, which is -5 . -24 + 1 = -23 -12 + 2 = -10 -8 + 3 = -5 That's it

OpenStudy (anonymous):

Any Q's? (:

OpenStudy (anonymous):

could you also find the answer by taking the list and going though and distributing them

OpenStudy (anonymous):

Yes… but you won't always have that option. Most exams these days are going more towards short answer vs. multiple choice like the EOC in FL.

OpenStudy (anonymous):

It isn't always about taking the easy way out. You must learn the process of how to do this at least. Plus it will take about the same amount of time.

OpenStudy (anonymous):

(4n - 3)(n + 2) 4ntimesn and 4ntimes 2 and -3timen

OpenStudy (anonymous):

That isn't all.

OpenStudy (anonymous):

i am taking summer school for math becouse i had a hard time during the school year

OpenStudy (anonymous):

Please learn the process with me. Distributing is primarily used to CHECK your answer at the end.

OpenStudy (anonymous):

Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, -8 and 3 4n^2 - 8n + 3n - 6

OpenStudy (anonymous):

Also.. don't be put down that you were put into summer school. It's actually better for you because you get to review the concepts at your own pace. Don't you dare be dishearted. Everyone is different and learn differently. Be confident! :)

OpenStudy (anonymous):

Is everything I've told you so far clear?

OpenStudy (anonymous):

Well… even though you're not here I'll just complete what I started:

OpenStudy (anonymous):

Step-4 : Add up the first 2 terms, pulling out like factors : 4n • (n-2) Add up the last 2 terms, pulling out common factors : 3 • (n-2) Step-5 : Add up the four terms of step 4 : (4n+3) • (n-2) Which is the desired factorization

OpenStudy (anonymous):

You were incorrect as it was the second one.